Circular No. 9022

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 33P/DANIEL Sometime between Jan. 11 and 30, comet 33P appears to have un-

dergone an outburst of ∼ 3 mag in total brightness (from mag ∼ 18 to ∼ 15), according to CCD astrometric observations contributed to the Minor Planet Center. CCD observations by K. Kadota (Ageo, Japan, 0.25-m f/5 reflector) and others indicate an additional brightening of ∼ 1 mag between Feb. 7 and 20. A standard power-law formula (cf. 2008/2009 Comet Handbook, p. 64) predicts little change in the comet’s brightness over this span (T = 2008 July 20.3 TT). Recent astrometry, revised orbital elements by B. G. Marsden, and an updated ephemeris appear on MPEC 2009-D52.

COMETS C/2008 U6, C/2008 U9–U16, C/2008 V1–V3 (SOHO) A revised orbit on MPEC 2009-B70 shows C/2008 U6 (cf. IAUC 9021)

to be a member of the Meyer group. Additional near-sun presumed comets have been found on SOHO website images, with their “discovery” observations tabulated below — all Kreutz sungrazers except for C/2008 V2 (Meyer group). C/2008 U9 was very diffuse (mag ∼ 8–8.5). C/2008 U10 and C/2008 U15 were faint (mag ∼ 8) and diffuse. C/2008 U11 was slightly diffuse; it appeared small (mag ∼ 7) in C3 images, with a short, diffuse tail in C2 images. C/2008 U12 and C/2008 V3 were very diffuse (mag ∼ 8). C/2008 U13 was stellar in appearance (mag ∼ 7.5) in C3 images, and quite condensed with no tail in C2 images. C/2008 U14 was slightly diffuse and elongated (mag ∼ 8). C/2008 U16 was small and condensed (mag ∼ 7.5– 8). C/2008 V1 was slightly diffuse; in C3 images, it was faint (mag ∼ 7.5). C/2008 V2 was stellar in appearance (mag ∼ 7).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 U9 Oct. 21.621 13^{h}39.^{m}7 −12^{o}03′ C2  HS 2009-B67
C/2008 U10 24.521 13 50.7 −13 06 C2 ZX 2009-B70
C/2008 U11  24.696 13 44.3 −13 52 C3/2 HS 2009-B70
C/2008 U12 26.329 13 57.7 −13 40 C2 BZ 2009-B70
C/2008 U13  26.988 13 53.6 −14 38 C3/2 RM 2009-B71
C/2008 U14  27.454 14 02.9 −14 10 C2  RK 2009-B71
C/2008 U15  28.413 14 06.6 −14 31 C2  JR 2009-B71
C/2008 U16  31.079 14 16.7 −15 23 C2  RM 2009-B71
C/2008 V1 Nov. 2.821 14 24.9 −16 55 C3/2 RM 2009-C46
C/2008 V2  4.246 14 31.5 −16 07 C2  RK 2009-C46
C/2008 V3  6.329 14 41.9 −17 25 C2  RM 2009-C46
2009 February 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green